# Atomistic analysis of nematic phase transition in 4-cyano-4-$n$-alkyl biphenyl liquid crystals: Sampling for the first-order phase transition and the free-energy decomposition

### v1

@@ I Introduction
To obtain a more precise molecular-level understanding beyond mean-field models, it is essential to analyze the free energy of mesogens inserted into both isotropic and nematic phases by varying the temperature across the NI phase transition of $n$CB, particularly through molecular dynamics (MD) simulations. However, this free-energy calculation remains challenging and has not been thoroughly elucidated, despite numerous MD simulations conducted on the NI phase transition of liquid crystal systems. [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]
@@
Figure 1: Temperature dependence of orientational order parameter $S$ for 5CB (a), 6CB (b), 7CB (c), and 8CB (d). The error bar at each temperature corresponds to the standard deviation. The vertical color bar indicates the NI transition temperature $T_{\mathrm{NI}}$.
In this study, we performed MD simulations of $n$CB system with $n=5$, 6, 7, and 8 using the united-atom (UA) model. We also applied the generalized replica-exchange method (gREM), designed to improve sampling of the states near the first-order phase transition by utilizing a non-Boltzmann wight. [30, 31, 32, 33, 34, 35, 36, 37, 38, 39] Note that gREM simulations of 5CB liquid crystals [40] and other applications of the replica-exchange method to coarse-grained liquid crystalline systems were conducted. [41, 42, 43, 44] The insertion free energy was subsequently evaluated using the energy representation (ER) theory for the states obtained through gREM, enabling the assessment of the thermodynamic stability of mesogens in both isotropic and nematic phases across varying temperatures. Within the ER framework, it is also possible to decompose the insertion free energy into the energetic and entropic contributions. We conduct free-energy decomposition and address the roles of these contributions.
@@ II Model and methods
The molecular model employed was the UA model of the AMBER force field, developed by Tiberio et al. to reproduce the experimental behavior near the NI phase transition temperature. [17] The system consists of $N=4000$ molecules; it was seen in Ref. 40 that the system size with $N=4000$ is enough for treatments of the NI transition. The simulation box is a cubic cell with periodic boundary conditions. All MD simulations in this study were performed using the Large-scale Atomic/Molecular Massively Parallel Simulator (LAMMPS) [45]. The time step was set to 2 fs. Temperature was maintained using Nosé–Hoover thermostat, while pressure was controlled by isotropic Nosé–Hoover barostat with a pressure of 1 atm.
For the isobaric version of gREM, the enthalpy $H(=U+pV)$ dependent non-Boltzmann weight $W_{\alpha}(H)$ $(\alpha=1,\cdots,M)$ is utilized. [34, 35, 36] Here, $U$, $p$, and $V$ represent the potential energy, pressure, and volume of the system, respectively. Here, $\alpha$ denotes the replica index and $M$ is the number of replicas. The weight $w_{\alpha}$ is connected with the effective potential through $w_{\alpha}=-k_{\mathrm{B}}\ln W_{\alpha}$, and is determined by inverse mapping of the effective temperature:
@@
The statistical temperature is defined as $T_{\mathrm{S}}(H)=[\partial S/\partial H]^{-1}$, where $S$ is the entropy. At replica $\alpha$, the statistical temperature $T_{\mathrm{S}}$ is evaluated by the most probable value of the enthalpy $H_{\alpha}^{*}$, with $T_{\alpha}(H_{\alpha}^{*})=T_{\mathrm{S}}(H_{\alpha}^{*})$. The simplest expression for $T_{\alpha}$ is the linear effective temperature, given by
@@
where $\lambda_{\alpha}$ and $\gamma$ are control parameters representing the intercept and slope of the straight line, respectively, at an arbitrary chosen enthalpy $H_{0}$. Note that $\gamma$ is chosen to be negative so that $T_{\alpha}(H)$ intersects $T_{\mathrm{S}}(H)$ at only one point. In practice, $\gamma$ is selected as
@@
where $\Delta_{\alpha,\alpha^{\prime}}=w_{\alpha^{\prime}}(H^{\prime})-w_{\alpha^{\prime}}(H)+w_{\alpha}(H)-w_{\alpha}(H^{\prime})$. In this study, the number of replicas was set to $M=17$. Other parameters, $T_{1}$, $T_{M}$, $\hat{H}_{1}$, $\hat{H}_{M}$, $\gamma$ are summarized in Table S1 of the supplementary material. The LAMMPS input files used in this study are open at the repository of Zenodo (see DATA AVAILABILITY statement).
Figure S1 of the supplementary material shows the distribution of the enthalpy normalized by the number of molecules, $H/N$, at each replica, for each replica for $n$CB ($n=5$, 6, 7, and 8). Figure S2 of the supplementary material displays the statistical temperature $T_{\mathrm{S}}$ of each temperature as a function of $H/N$, with the linear effective temperature, as defined in Eq. (2), represented by straight lines. Henceforth, the statistical temperature $T_{\mathrm{S}}$ is referred to simply as $T$.
@@
where the first term represents the average sum of solute-solvent interactions in the target system with $\lambda=1$, and the second term accounts for the free-energy penalty due to the insertion of the solute molecule in Eq. (7). Furthermore, in the ER theory, the second term is approximated by a density functional form $\mathcal{F}$ using $\rho_{0}(\varepsilon)$ with $\lambda=0$, representing the reference solvent system, and $\rho(\varepsilon)$ with $\lambda=1$, representing the target system, as expressed in Eq. (8). In other words, ER theory approximates the insertion free energy from the energy distribution functions of the initial ($\lambda=0$) and final ($\lambda=1$) states, thereby reducing the computational cost by avoiding MD simulations of intermediate states. It has also been reported that the error due to the use of density functional approximation is no greater than the inherent error in the force field. [51] In Eq. (8), the first term is the average interaction energy of the solute with the solvent in the system of interest. The second term corresponds to the free-energy penalty due to the reorganization of the solvent structures with the insertion of the solute. When the solvent degrees of freedom are more restricted by the solute, the second term is more positive. In the following, the first and second terms of Eq. (8) are called energetic and entropic terms, respectively.
@@ III.1 Orientational order
where $\theta_{i}$ represents the orientational angle between the $i$-th molecular axis $\bm{u}_{i}$ and the director axis $\bm{n}$. Furthermore, $P_{2}(x)=(3x^{2}-1)/2$ is the second-order Legendre polynomial. In this study, the molecular axis was defined as the CN bond. The second-rank order parameter tensor,
@@
is employed to describe the orientational order of biaxial nematic liquid crystals. Here, $\bm{I}$ denotes the identity matrix. Due to the traceless property of $\bm{Q}$, three eigenvalues, $\lambda_{-}$, $\lambda_{0}$, and $\lambda_{+}$ ($\lambda_{-}<\lambda_{0}<\lambda_{+}$), satisfies the condition $\lambda_{-}+\lambda_{0}+\lambda_{+}=0$. The maximum eigenvalue $\lambda_{+}$ coincides with the orientational order parameter $S$, and the corresponding eigenvector represents the director $\bm{n}$. To prevent the order parameter from taking large positive values, particularly in the isotropic phase, $S=-2\lambda_{0}$ was calculated in practice. [52]
The temperature dependence of the ensemble average of S is plotted in Fig. 1. In all liquid crystal system, $S$ exhibits a value near 0 at high temperatures, indicating the isotropic phase. As the temperature decreases a discontinuous behavior is observed at a certain temperature, and $S$ reaches a value around 0.4, signaling a phase transition to the nematic phase. The highest temperature at which $S$ reaches approximately 0.4 was defined as the NI phase transition temperature, $T_{\mathrm{NI}}$. The $S$ value gradually increases as the temperature decreases further, indicating a strengthening of the nematic phase’s orientational order. Although 8CB of the same UA model is known to undergo a smectic phase transition upon further cooling, [17, 19] this study specifically targets the NI phase transition and does not address the smectic-nematic phase transition. Simulated $T_{\mathrm{NI}}$ values for $n$CB ($n=5$, 6, 7, and 8) are 308.7 K (replica 8), 300.0 K (replica 9), 324.3 K (replica 12), and 318.9 K (replica 8), respectively. The results demonstrate that $n$CB exhibits odd-even effects as the number of carbons in the alkyl chain increases, consistent with findings from other MD simulations. [10, 12, 14, 16, 17]
@@ III.2 Comparison with Maier–Saupe model
where $\Gamma_{0}$ characterizes the strength of intermolecular interactions. Equation (11) represents the Boltzmann distribution of $V(\cos\theta,S)$, thereby classifying the MS model as an attractive model.
@@
The first term can be regarded as the ensemble average of the intermolecular potential, $\langle V(\cos\theta,S)/k_{\mathrm{B}}T\rangle/2=\Gamma_{0}S^{2}/2$, where the factor $1/2$ compensates for the double-counting of interactions. The second term corresponds to the Kullback–Leibler divergence, which quantifies the difference between the distribution $P(\cos\theta)$ and the uniform distribution characterizing the isotropic phase, $1/4\pi$. By using $\Gamma_{0}$ as the control parameter, the orientation order parameter $S$ can be evaluated numerically in the self-consistent manner.
At the NI phase transition, the parameter $\Gamma_{0}\approx 4.54$ yields an orientation order parameter value of $S\approx 0.4295$, which coexists with the trivial solution of $S=0$. Using the corresponding NI phase transition temperature $T_{\mathrm{NI}}$, $S$ as a function of $T/T_{\mathrm{NI}}$ for the MS model is plotted in Fig. 2(a). A comparison with MD simulations for $n$CB ($n=5$, 6, 7, and 8) is also presented, showing overall consistency between MD simulations and MS model. Furthermore, the orientation angle distributions, $P(\cos\theta)$, at $T_{\mathrm{NI}}$ for the MS model is illustrated in Fig. 2(b). The result presents the agreement with MD simulation results for $n$CB, particularly highlighting a significant proportion of molecules aligned parallel to the directer. Slight deviations between MD simulations and the MS model are observed in Fig. 2, which are attributed to the value of $S$ at $T_{\mathrm{NI}}$ being less than the mean-field value of $S=0.4295$.
@@ III.4 Energy representation free-energy analysis
Finally, we analyze the thermodynamic stability of the mesogen using the gREM sampling near the NI phase transition temperature. Specifically, we calculated the temperature dependence of the insertion free energy in the $n$CB system based on the ER thory (see Eq. (8)). The results are shown in Fig. 4. Note that $\rho(\varepsilon)$ and $\rho_{0}(\varepsilon)$ were evaluated through MD simulations of systems with $N=4000$ and $N=3999$ molecules, respectively. The results of $\rho(\varepsilon)$ and $\rho_{0}(\varepsilon)$ of the replica at $T_{\mathrm{NI}}$ and its neighboring replicas (isotropic and nematic phases) are presented in Fig. S3 and Fig. S4 of the supplementary material, respectively.
@@ IV Conclusions
The supplementary material includes parameters of gREM simulations (Table S1); probability distribution of enthalpy of replicas in gREM simulations (Fig. S1); statistical temperature as a function of enthalpy in gREM simulations (Fig. S2); density profiles of intermolecular interaction energy for ER calculations, $\rho(\varepsilon)$ and $\rho_{0}(\varepsilon)$ (Fig. S3 and Fig. S4); and temperature dependence of LJ and electrostatic energies (Fig. S5).

### v2

@@ I Introduction
To obtain a more precise molecular-level understanding beyond mean-field models, it is essential to analyze the free energy of mesogens inserted into both isotropic and nematic phases by varying the temperature across the NI phase transition of $n$CB, particularly through molecular dynamics (MD) simulations. However, this free-energy calculation remains challenging and has not been thoroughly elucidated, despite numerous MD simulations conducted on the NI phase transition of liquid crystal systems. [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]
@@
Figure 1: Temperature dependence of orientational order parameter $S$ for 5CB (a), 6CB (b), 7CB (c), and 8CB (d). The molecular axis is defined using the CN bond, represented by the purple vector in Fig. S1 of the supplementary material. The error bar at each temperature corresponds to the standard deviation. The vertical color bar indicates the NI transition temperature $T_{\mathrm{NI}}$.
In this study, we performed MD simulations of $n$CB system with $n=5$, 6, 7, and 8 using the united-atom (UA) model. We also applied the generalized replica-exchange method (gREM), designed to improve sampling of the states near the first-order phase transition by utilizing a non-Boltzmann weight. [31, 32, 33, 34, 35, 36, 37, 38, 39, 40] Note that gREM simulations of 5CB liquid crystals [41] and other applications of the replica-exchange method to coarse-grained liquid crystalline systems were conducted. [42, 43, 44, 45] The insertion free energy was subsequently evaluated using the energy representation (ER) theory for the states obtained through gREM, enabling the assessment of the thermodynamic stability of mesogens in both isotropic and nematic phases across varying temperatures. Within the ER framework, it is also possible to decompose the insertion free energy into the energetic and entropic contributions. We conduct free-energy decomposition and address the roles of these contributions.
@@ II Model and methods
The molecular model employed was the UA model of the AMBER force field, developed by Tiberio et al. to reproduce the experimental behavior near the NI phase transition temperature. [17] The molecular structures of $n$CB ($n=5$, 6, 7, and 8) are illustrated in Fig. S1 of the supplementary material. The system consists of $N=4000$ molecules; it was seen in Ref. 41 that the system size with $N=4000$ is enough for treatments of the NI transition. The simulation box is a cubic cell with periodic boundary conditions. All MD simulations in this study were performed using the Large-scale Atomic/Molecular Massively Parallel Simulator (LAMMPS) [46]. The time step was set to 2 fs. The total simulation times of gREM were 360 ns for 5CB and 6CB and 450 ns for 7CB and 8CB, respectively, starting from an initial configuration with a random structure. Temperature was maintained using Nosé–Hoover thermostat, while pressure was controlled by isotropic Nosé–Hoover barostat with a pressure of 1 atm. For the isobaric version of gREM, the enthalpy $H(=U+pV)$ dependent non-Boltzmann weight $W_{\alpha}(H)$ $(\alpha=1,\cdots,M)$ is utilized. [35, 36, 37] Here, $U$, $p$, and $V$ represent the potential energy, pressure, and volume of the system, respectively. In addition, $\alpha$ denotes the replica index and $M$ is the number of replicas. The weight $w_{\alpha}$ is connected with the effective potential through $w_{\alpha}=-k_{\mathrm{B}}\ln W_{\alpha}$, and is determined by inverse mapping of the effective temperature:
@@
The statistical temperature is defined as $T_{\mathcal{S}}(H)=[\partial\mathcal{S}/\partial H]^{-1}$, where $\mathcal{S}$ is the entropy. At replica $\alpha$, the statistical temperature $T_{\mathcal{S}}$ is evaluated by the most probable value of the enthalpy $H_{\alpha}^{*}$, with $T_{\alpha}(H_{\alpha}^{*})=T_{\mathcal{S}}(H_{\alpha}^{*})$. The simplest expression for $T_{\alpha}$ is the linear effective temperature, given by
@@
where $\lambda_{\alpha}$ and $\gamma$ are control parameters representing the intercept and slope of the straight line, respectively, at an arbitrary chosen enthalpy $H_{0}$. Note that $\gamma$ is chosen to be negative so that $T_{\alpha}(H)$ intersects $T_{\mathcal{S}}(H)$ at only one point. In practice, $\gamma$ is selected as
@@
where $\Delta_{\alpha,\alpha^{\prime}}=w_{\alpha^{\prime}}(H^{\prime})-w_{\alpha^{\prime}}(H)+w_{\alpha}(H)-w_{\alpha}(H^{\prime})$. In this study, the number of replicas was set to $M=17$. Other parameters, $T_{1}$, $T_{M}$, $\hat{H}_{1}$, $\hat{H}_{M}$, $\gamma$ are summarized in Table S1 of the supplementary material. The LAMMPS input files used in this study and 1 ns trajectory files are openly available at the repository of Zenodo (see DATA AVAILABILITY statement).
Figure S2 of the supplementary material shows the distribution of the enthalpy normalized by the number of molecules, $H/N$, at each replica, for each replica for $n$CB ($n=5$, 6, 7, and 8). These results were calculated from the final 36 ns (5CB and 6CB) and 108 ns (7CB and 8CB) trajectories in the gREM calculation. The findings demonstrate sufficient overlap of enthalpy distributions between replicas. However, the distributions are broad in certain regions of enthalpy between specific replicas, which can be attributed to enthalpy gaps caused by the NI phase transition. Despite this, we confirmed that the replica exchange acceptance ratio remains maintained, as shown in Fig. S3 of the supplementary material. To enhance the replica exchange acceptance ratio, particularly for larger system sizes, it is necessary to increase the number of replicas $M$, significantly raising computational demands. Figure S4 of the supplementary material displays the statistical temperature $T_{\mathcal{S}}$ of each temperature as a function of $H/N$, with the linear effective temperature, as defined in Eq. (2), represented by straight lines. Henceforth, the statistical temperature $T_{\mathcal{S}}$ is referred to simply as $T$.
@@
where the first term represents the average sum of solute-solvent interactions in the target system with $\lambda=1$, and the second term accounts for the free-energy penalty due to the insertion of the solute molecule in Eq. (7). Furthermore, in the ER theory, the second term is approximated by a density functional form $\mathcal{F}$ using $\rho_{0}(\varepsilon)$ with $\lambda=0$, representing the reference solvent system, and $\rho(\varepsilon)$ with $\lambda=1$, representing the target system, as expressed in Eq. (8). In other words, ER theory approximates the insertion free energy from the energy distribution functions of the initial ($\lambda=0$) and final ($\lambda=1$) states, thereby reducing the computational cost by avoiding MD simulations of intermediate states. It has also been reported that the error due to the use of density functional approximation is no greater than the inherent error in the force field. [52] Specifically, the ER method was used to calculate the solvation free energy of amino acid analogs in a pure-water solvent, showing a typical discrepancy of less than 1 kcal/mol compared to the numerically exact values obtained using the Bennett acceptance ratio (BAR) method. [52] In Eq. (8), the first term is the average interaction energy of the solute with the solvent in the system of interest. The second term corresponds to the free-energy penalty due to the reorganization of the solvent structures with the insertion of the solute. When the solvent degrees of freedom are more restricted by the solute, the second term is more positive. In the following, the first and second terms of Eq. (8) are called energetic and entropic terms, respectively.
@@ III.1 Orientational order
where $\theta_{i}$ represents the orientational angle between the $i$-th molecular axis $\bm{u}_{i}$ and the director $\bm{n}$. Furthermore, $P_{2}(x)=(3x^{2}-1)/2$ is the second-order Legendre polynomial.
In MD simulations, the direction of the director $\bm{n}$ cannot be predetermined. Alternatively, the second-rank order parameter tensor,
@@
is employed to analyze the orientational order of biaxial nematic liquid crystals. Here, $\bm{I}$ denotes the identity matrix. Due to the traceless property of $\bm{Q}$, three eigenvalues, $\lambda_{-}$, $\lambda_{0}$, and $\lambda_{+}$ ($\lambda_{-}<\lambda_{0}<\lambda_{+}$), satisfies the condition $\lambda_{-}+\lambda_{0}+\lambda_{+}=0$. The orientational order parameter $S$ can be evaluated by the maximum eigenvalue $\lambda_{+}$ as $S=\lambda_{+}$, and the corresponding eigenvector represents the director $\bm{n}$. However, this definition results in positive values for $S$ even in the isotropic phase due to the traceless property of $Q$. In this study, following the approach of Eppenga and Frenkel, $S$ was calculated by $-2\lambda_{0}$, considering $\lambda_{0}\approx\lambda_{-}=-\lambda_{+}/2$ in the nematic phase. [53] This definition ensures that $S$ fluctuates around zero in the isotropic phase, indicating the absence of orientational order.
The temperature dependence of the ensemble average of $S$ is plotted in Fig. 1. In this study, the molecular axis was defined as the CN bond, represented by the purple vector in Fig. S1 of the supplementary material. Here, $S$ was calculated from the final 36 ns (5CB and 6CB) and 108 ns (7CB and 8CB) trajectories in the gREM simulation, consistent with the calculation of the enthalpy distribution shown in Fig. S2 of the supplementary material. In all liquid crystal systems, $S$ exhibits a value near 0 at high temperatures, indicating the isotropic phase. As the temperature decreases, a discontinuous behavior is observed at a certain temperature, and $S$ reaches a value around 0.4, signaling a phase transition to the nematic phase. The highest temperature at which $S$ reaches approximately 0.4 was defined as the NI phase transition temperature, $T_{\mathrm{NI}}$. The $S$ value gradually increases as the temperature decreases further, indicating a strengthening of the nematic phase’s orientational order. Although 8CB of the same UA model is known to undergo a smectic phase transition upon further cooling, [17, 19] this study specifically targets the NI phase transition and does not address the smectic-nematic phase transition. Simulated $T_{\mathrm{NI}}$ values for $n$CB ($n=5$, 6, 7, and 8) are 308.7 K (replica 8), 300.0 K (replica 9), 324.3 K (replica 12), and 318.9 K (replica 8), respectively. The results demonstrate that $n$CB exhibits odd-even effects as the number of carbons in the alkyl chain increases, consistent with findings from other MD simulations. [10, 12, 14, 16, 17] We also calculated the orientational order parameter $S$ using an alternative definition of the molecular axis, defined by the line connecting N and C of the phenyl ring, represented by the red vector in Fig. S1 of the supplementary material. This definition was examined in a previous study. [30] The temperature dependence of $S$, shown in Fig. S5 of the supplementary material, demonstrates that $S$ remains unaffected by the choice of the molecular axis.
@@ III.2 Comparison with Maier–Saupe model
where $\Gamma_{0}$ characterizes the strength of intermolecular interactions. Equation (11) represents the Boltzmann distribution of $V(\cos\theta,S)$, thereby classifying the MS model as an attractive model. In this context, the MS model considers the combined effects of volume fraction and temperature on the NI phase transition, making it more suitable for describing thermotropic liquid crystals compared to the Onsager model.
@@
The first term can be regarded as the ensemble average of the intermolecular potential, $\langle V(\cos\theta,S)/k_{\mathrm{B}}T\rangle/2=\Gamma_{0}S^{2}/2$, where the factor $1/2$ compensates for the double-counting of interactions. The second term corresponds to the Kullback–Leibler divergence, which quantifies the difference between the distribution $P(\cos\theta)$ and the uniform distribution characterizing the isotropic phase, $1/4\pi$. By using $\Gamma_{0}$ as the control parameter, the orientational order parameter $S$ can be evaluated numerically in the self-consistent manner based on Eq. (14).
At the NI phase transition, the parameter $\Gamma_{0}\approx 4.54$ yields an orientation order parameter value of $S\approx 0.4295$, which coexists with the trivial solution of $S=0$. Using the corresponding NI phase transition temperature $T_{\mathrm{NI}}$, $S$ as a function of $T/T_{\mathrm{NI}}$ for the MS model is plotted in Fig. 2(a). A comparison with MD simulations for $n$CB ($n=5$, 6, 7, and 8) is also presented, showing overall consistency between MD simulations and MS model. Furthermore, the orientation angle distributions, $P(\cos\theta)$, at $T_{\mathrm{NI}}$ for the MS model is illustrated in Fig. 2(b). The result presents the agreement with MD simulation results for $n$CB, particularly highlighting a significant proportion of molecules aligned parallel to the director. Slight deviations between MD simulations and the MS model are observed in Fig. 2, which are attributed to the value of $S$ at $T_{\mathrm{NI}}$ being less than the mean-field value of $S=0.4295$.
@@ III.4 Energy representation free-energy analysis
We analyze the thermodynamic stability of the mesogen using the gREM sampling near the NI phase transition temperature. Specifically, we calculated the temperature dependence of the insertion free energy, $\Delta\mu$, in the $n$CB system based on the ER thory (see Eq. (8)). The results are shown in Fig. 4. Note that $\rho(\varepsilon)$ and $\rho_{0}(\varepsilon)$ were evaluated through MD simulations of systems with $N=4000$ and $N=3999$ molecules, respectively. The results of $\rho(\varepsilon)$ and $\rho_{0}(\varepsilon)$ of the replica at $T_{\mathrm{NI}}$ and its neighboring replicas (isotropic and nematic phases) are presented in Fig. S6 and Fig. S7 of the supplementary material, respectively.
@@
Finally, we discuss the odd-even effect based on our free-energy calculation of Fig. 5. For odd-numbered alkyl chains, the terminal alkyl group aligns parallel to the molecular axis, while for even-numbered alkyl chains, the angle to the molecular axis tends to be larger. This results in reduced molecular anisotropy in the even-numbered compounds, leading to a lower NI phase transition temperature. We analyzed the energy gap and entropy gap across the NI transition temperature, denoted as $\Delta\langle u\rangle$ and $\Delta(\Delta\mu-\langle u\rangle)$. As illustrated in Fig. S9 of the supplementary material, the results indicate that odd-numbered compounds exhibit more stability of the nematic ordering with a significantly large energy loss, accompanied by a large entropy loss. This suggests that the nematic phase in odd-numbered compounds is stabilized by a strong molecular anisotropy. Conversely, the even-numbered compounds show smaller energy and entropy gaps, reflecting their reduced anisotropy and lower transition temperature to the nematic phase.
@@ IV Conclusions
The supplementary material includes parameters of gREM simulations (Table S1); molecular structures of $n$CB (Fig. S1); probability distribution of enthalpy of replicas in gREM simulations (Fig. S2); acceptance ratios of replica exchanges (Fig. S3); statistical temperature as a function of enthalpy in gREM simulations (Fig. S4); orientational order parameter $S$ using another molecular axis (Fig. S5); density profiles of intermolecular interaction energy for ER calculations, $\rho(\varepsilon)$ and $\rho_{0}(\varepsilon)$ (Fig. S6 and Fig. S7); and temperature dependence of LJ and electrostatic energies (Fig. S8); dependence of energy gap $\Delta\langle u\rangle$ and entropy gap $\Delta(\Delta\mu-\langle u\rangle)$ on the number of carbons in the alkyl chain (Fig. S9).